Hello what is the hypotenuse of :
Leg a : 18cm
Leg b : 24cm

Answer by sheldonbbtrocks(115)

a^2+b^2=c^2 Where a and b are the legs of the triangle and c is the hypoteneuse:
24^2+18^2=c^2
576+324=c^2
900=c^2
Square root both sides: 300=c
Also note that this triangle: 24,18,30 is called a Pythagorean Triple. It is a fairly common one!
Hope this helps!